A tire at 32 PSI at 37oC is left outside and the temperature drops to 10oC overnight. What is the pressure in the tire at the new temperature? 1 atm = 14.7 lb/in2
Answer #1

Answer:

From Ideal Gas equation, PV=nRT

Given pressure =32 PSI at temperature=37°C

and at temperature=10°C waht is pressure.

Here we have to use the formula P/T= constant. (Since R = gas constant and volume is fixed=constant)

P1/T1=P2/T2

Given P1=32 PSI=2.177 atm. (Since 1 PSI=0.068 atm or 1 atm=14.7 PSI or lb/in2)

T1=37°C and T2=10°C

P2=? then P2=P1T2/T1=(2.177 atm x 10°C)/37°C

P2=0.588 atm or 8.64 PSI.